OpenStudy (anonymous):

Solve for x: (x-4)2 =2 the work: first simplify (x-4)2 which = 2x-8 (2 times x + 2 times -4) 2x-8=2 then you have to bring the -8 over by the other side by adding 8 to both sides. Whatever you do on one side you have to do on the other to balance the equation. 2x-8+8 = 2+8 2x = 10 Divide both sides by 2 to get x by itself. 2x/2 = 10/2 x= 5 BUT..... which one is it? x = 4 ± 2 sq root 2 x = 4 ± sq root 13 x = 4 ± sq root 2 x = 4 ± 2 sq root 13

OpenStudy (anonymous):

why you are making the simple problem too difficult....

OpenStudy (anonymous):

:/ ouch

OpenStudy (anonymous):

I was solving it is all

OpenStudy (anonymous):

( x - 4 )2 = 2 ( x - 4 ) = 1 dividing both sides by 2 x = 5 adding 5 to both sides

OpenStudy (anonymous):

It still doesn't help me

OpenStudy (anonymous):

no offense

OpenStudy (anonymous):

I think you are write the question something is missing. Do you mean (x-4)^2=2?